What is the next number in the sequence 2, 1, 3, 4, 7, 11, …?
Answer by Targetweek(62)

well 2 + 1 = 3
then 1 + 3 = 4
then 3 + 4 = 7
then 4 + 7 = 11
then 7 + 11 = 18
The answer is 18